slope of a tan line to a curve at P (3,3) is 4. Use the equation of this tangent to find its value for x=9.
The answer I got was p=81 , I know that the equation of the tangemt line is y=4x-9, but I am lost after that part.
but my choices are 27,31,29,22,32 or 24 as the right answer. so it can not be p=81

Answer by Alan3354(69443)

slope of a tan line to a curve at P (3,3) is 4. Use the equation of this tangent to find its value for x=9.
------------------
y-y1 = m*(x-x1)
y-3 = 4(x-3)
y = 4x - 9
--------
At x = 9:
y = 4*9 - 9
y = 27
You got the hard part right.